# Building spatial world models from sparse transitional episodic memories

### v1

@@ 4 Results
4.1 Simulating Spatial Environments
We first assessed ESWM on its core ability to predict missing elements of unseen transitions given a set of memories from an environment and a partially observable query. We found a stark difference between the model types in Open Arena setting (Fig.2a). The transformer model excelled at learning this task, reaching accuracies of 82.9%, 95.8%, and 85.4% on source state, action, and end state prediction tasks respectively. In contrast, LSTM and Mamba models struggled to reach above chance on this task for both training and testing. This suggests that the attention mechanism in Transformer architectures—reminiscent of classic content-addressable memory models (29)—plays a key role in learning generalizable world models from episodic memories. However, a closely related approach by Coda-Forno et al. (8), which relies on a key-value memory table and hand-engineered K-nearest neighbor lookup, fails to predict unseen queries (Fig. A8). This highlights the importance of using learnable, deep attention-based mechanisms for flexible inference.
@@ 4.1 Simulating Spatial Environments
In the second scenario, the Random Wall, we found that all models were capable of learning the task (Fig.2b). Although, there was a substantial gap between transformer-based models and other models, even with equal number of parameters (LSTM-4L, Transformer-2L, and Mamba-5L). This showed that ESWM learned a general strategy for modeling an environment from sparse and disjoint episodic memories of transitions within that environment that it could then generalize to any environment with different structures.
4.2 Spatial Maps Emerge Robustly From the Latent Space
To simulate an unseen transition, ESWM must integrate disjoint memory fragments into a coherent internal map that reflects both observed states and obstacles. Inspired by low-dimensional, environment-like neural manifolds in rodents’ MTL (38), we applied ISOMAP (49) (Extended Methods A.5) to ESWM’s latent space. The resulting manifolds revealed that ESWM consistently organizes its internal representations to mirror the spatial layout of each specific environment.
In obstacle-free environments, the manifold adopts a smooth, saddle-like shape (Fig.3a). When obstacles are introduced, they create localized discontinuities that correspond precisely to blocked regions in physical space (Fig.3b). This structure is remarkably robust and persists under out-of-distribution obstacle shapes (Fig. A1, A2), partial observation of environments (Fig. A3), across architectures (LSTM, Transformer, Mamba; Figs.A1A2), and throughout all transformer layers, which become progressively more continuous with depth (Fig.A5). All prediction tasks (start-state, action, end-state) yield similar maps (Fig.A3a). Quantitatively, path lengths in latent and physical spaces correlate strongly ($R^{2}=0.89$, $N_{\text{path}}=1500$, $N_{\text{env}}=75$; Fig.A7b), and a linear classifier on ESWM activations can accurately discriminate which of two states is farther from a reference state with $93.42\pm 0.006\%$ accuracy (chance = 50%, $N_{\text{seed}}=10$; Extended Methods A.6).
Notably, such structured spatial representation do not emerge in episodic agents trained for navigation, such as Episodic Planning Network (EPN; (42)), despite being trained with meta-RL on the same environments and data as ESWM (Fig.A1, A2; Appendix D).
(a)(b)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c)(d)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 4.2 Spatial Maps Emerge Robustly From the Latent Space
Figure 3: Spatial map emerges in ESWM’s latent space. a), b) ISOMAP projections of the 2-layer transformer model’s first layer’s activations for action prediction. Different columns are the same projection from different viewing angles. a) Spatial map in the absence of obstacles and b) in the presence of obstacles (a straight wall). c), d) ISOMAP projections of the 14-layer transformer model’s seventh layer’s activations. c) From left to right: A sample room with two disjoint regions whose shapes match boundaries’ shape; ESWM’s latent space when a memory bank observing either the top or bottom region is given as input; ESWM’s latent space when a memory bank observing both regions is given as input. d) From left to right: A sample room containing two disjoint regions whose shapes give no cues about their relative position; ESWM’s latent space when a memory bank of the room is given as input; Updated room with a new wall, the regions remain disjoint; ESWM’s latent space when a memory bank of the updated room is given as input. See Fig.A1, A2, A3, A4, A5 for more examples.
4.3 Handling Disjoint Memory Clusters
During Random Wall training, memory banks can contain disjoint memory clusters, with no paths between observations in distinct clusters. This leads to uncertainty in spatial relationships as path integration provides no avail. In rare cases, the model is asked to solve queries that bridge disconnected clusters. This raises two questions: (1) Does the model infer the relative positions of disjoint memory clusters (e.g., which part of the room each cluster maps onto)? (2) Which cues enable these inferences?
4.3.1 Inference from Environmental Structure
We observe that when memory clusters are disjoint, the model leverages the underlying structure of the room to infer their relative positions. For example, the model aligns memory clusters with the room’s boundary when their shape matches. This allows the model to piece together disjoint memory clusters, like solving a puzzle, by identifying the configuration that satisfies the invariant structural constraints. Evidence for this behaviour is twofold. First, in the model’s latent space, the disjoint clusters are stitched together, rather than overlapping or disconnecting, to form a smooth surface reflecting the room’s structure (Fig. 3c). Second, the model’s end state prediction distributions, starting from the boundary states of one cluster, are biased towards the boundary states of the inferred neighbouring cluster, instead of exclusively outputting the “I don’t know” option.
4.3.2 Generalization from Observations of Obstacles
In cases where disjoint memory clusters’ shapes do not provide clear cues about their relative locations, the model relies on observations of obstacles to deduce their positions. For instance, given two disjoint clusters of memories, the model initially maps them to different locations in its latent space and outputs predictions exclusively reflecting uncertainty (“I don’t know”). However, if both clusters observe obstacles, even though the clusters remain disjoint, the model applies its knowledge of common wall shapes (e.g, walls are usually straight) to make inferences about the relative locations of the clusters. In the latent space, the two previously disconnected clusters are connected and aligned along the inferred wall shape to form a continuous representation (Fig. 3d, A4). The model’s predictions are also biased towards the inferred neighbours. Moreover, a single observation of obstacles from both clusters is sufficient for the model to infer the shape of the wall, and therefore infer the relative positions of the clusters (Fig. A4). These results demonstrate that ESWM can leverage the invariant structure of the room and combine observations of obstacles with prior knowledge of wall shapes to construct a coherent spatial representation.
4.4 ESMW Integrates Overlapping Memories
How does ESWM construct its internal spatial map? We find that ESWM’s prediction uncertainty—measured by entropy—increases with the length of the integration path needed to solve a query. That is, when ESWM must stitch together more episodic memories via overlapping states, uncertainty rises (Fig.4a). Introducing shortcut memories that reduce this path length significantly alters the prediction distribution (independent two-sample t-test; $n=5000$, $p<0.001$; Fig.4b). These results suggest that ESWM builds coherent spatial representations by integrating overlapping episodes. Consistent with this, performance improves with denser memory banks, even though they are out-of-distribution (Fig. 4c).
(a)(b)
@@ 4.4 ESMW Integrates Overlapping Memories
Figure 4: ESWM integrates memories. a) X-axis represents the shortest path length between the source and end states in the query, with edges corresponding to episodic memories in the memory bank, and the Y-axis is the entropy of ESWM’s prediction probabilities (n=5000). b) KL divergence between ESWM prediction distributions before and after adding an episodic memory to the memory bank. An informative episodic memory shortens the integration path required for the model to solve the prediction task, while a non-informative episodic memory does not alter the shortest integration path (n=5000). c) Prediction accuracies for unseen transistions as ESWM receive larger, out-of-distribution, memory banks (n=2000). Transformer -14L is used for a) and b) while 4L is used for c).
@@ 4.5 Exploration in Novel Environments
We designed a simple exploration algorithm based on ESWM’s “I don’t know” predictions (Alg.2). Starting with an empty memory bank, the agent evaluates all possible actions at each step by querying ESWM. It selects the action with the highest predicted uncertainty (i.e., highest “I don’t know” probability) and adds the resulting transition to memory. When ESWM is confident about all immediate actions, the agent performs multi-step lookahead by unrolling the model to identify frontier states with low prediction confidence (60). A prediction is deemed unconfident if ESWM either outputs “I don’t know” or assigns $\leq 80\%$ probability to the predicted outcome.
Interestingly, the agent adopts an efficient zig-zag exploration strategy, only revisiting states when separated by long paths. In the Random Wall environment, ESWM explores 16.8% more unique states than EPN (one-way ANOVA; $n=1000$, $F=2803.63$, $p<0.05$; Tukey’s HSD post hoc test: ESWM > EPN, $p<10^{-3}$), and achieves 96.48% of the performance of an oracle agent with full obstacle knowledge (Fig. A6). When exploration continues until memory is saturated, ESWM visits on average 90% of all unique states. Moreover, the collected memory banks allow ESWM to predict transitions with 96.71% of the accuracy of those achieved using procedurally generated memory banks. Together, these results highlight ESWM’s ability to support data-efficient, near-oracle exploration and mapping in novel environments.
@@
Figure 5: Exploration, navigation, and adaptability. a) Comparison of exploration strategies based on the number of unique states visited over 15 time steps in Random Wall. The optimal agent explores along the path found by the Traveling Salesman Algorithm over known free space. N=1000. b) Comparison of navigation success rate and path optimality over path lengths between EPN and ESWM, n=2400. c) Comparison of navigation success rates with increasing number of unexpected obstacles. The baseline agent is trained on the original environment and tested after structural changes, while ESWM navigates with memory banks from the original environment and needs to autonomously adapt to changes (n=100). Random Wall Experiments include 19 locations.
@@ 4.6 Navigating with Minimal Spanning Episodic Memories
Internal world models offer the ability to simulate long trajectories — imagining sequences of actions and observations beyond immediate sensory input— towards a certain goal before executing them. We demonstrate that memory-bank-equipped ESWM can be used for such long trajectory planning in obstacle-rich Random Wall. We frame the planning problem as follows: Given a memory bank of an environment and a list of possible actions, an agent must find a sequence of actions to go from any source state $S_{\text{source}}$ to any goal state $S_{\text{goal}}$. The agent operates without global information, such as its current location, the positions of obstacles, or the coordinates of the goal. Instead, it relies solely on its current (egocentric) observation, goal state, and fragmented spanning episodic experiences.
4.6.1 Planning by Imagination
@@
We devise a simple strategy relying on ESWM’s ability to infer the end state of any transition given a memory bank, the source state, and the action. Initialized at a state, an agent can imagine multiple steps into the future by querying ESWM on the consequences of all applicable actions, bootstrapping on ESWM’s own predicted end states, and tracking visited states until the goal state is reached (Alg.1). Then the agent can act out the imagined sequence of actions in reality. This can be viewed as Dijkstra’s algorithm applied to a graph whose dynamics can be retrieved through ESWM.
@@
Using this strategy, an ESWM-based navigational agent can navigate between arbitrary states in Random Wall with 96.8% success rate and 99.2% path optimality, outperforming EPN (42) by 18% and 21% respectively (Fig.5b) (one-sided independent t-test; n=2400: ESWM>EPN on optimality; p<$10^{-3}$). Navigation is successful if the agent reaches the target within 15 time steps, and path optimality is defined as the ratio between the shortest and actual path length.
4.6.2 Planning by Guided Imagination
While the previous strategy enables planning, it uniformly searches through the space, unbiasedly imagining the consequences of all actions, even if some move further from the goal. In contrast, animals with internal maps encoding geometric constraints plan biasedly, prioritizing actions that shorten the geodesic distance (i.e, the shortest path length) to the goal. We show that the spatial maps in ESWM’s latent space, which accurately encode obstacles and geodesic distance between states (Fig.A7b), can be used for guided imagination towards the goal.
The previous strategy is akin to Dijkstra’s algorithm, or A* with a null heuristic function (24; 15). The key to efficient planning, which prunes undesirable actions early, then, is a heuristic function containing rich information on the geodesic distances between states in the environment. Inspired by ISOMAP (49), we construct such a heuristic function $h_{\text{latent}}$ as follows (See Alg.4 for pseudocode): 1) Query the model to solve the end state prediction task for all possible $(s,a)$ pairs, collecting activations. 2) Construct a graph $G_{\text{latent}}$ where nodes are activations and edges connect nodes within radius $R_{\text{latent}}$ (Appendix 4). Edges are weighted by the cosine distance between their two endpoints (Fig.A7b). Intuitively, each node on $G_{\text{latent}}$ is the activation of predicting the outcome of action $a$ starting at state $s$. 3) Compute pairwise shortest paths on this graph to create a geodesic distance table $h_{\text{latent}}$.
$h_{\text{latent}}$ captures ESWM’s latent spatial map into a look-up table containing information on which action takes you closer to the goal state at every step of imagination. A* guided imagination with $h_{\text{latent}}$, improves planning efficiency by 57% (Fig.A7a), as measured by the number of states visited, over Dijkstra’s while maintaining path optimality (Fig.A7c). Furthermore, greedy policy based on $h_{\text{latent}}$ achieves 70% success in navigating around obstacles (Fig.A7a). These results suggest that structured latent spaces in world models can serve as efficient heuristics for planning.

### v2

@@ 4 Results
4.1 Modeling Spatial Environments from Episodic Memories
We first assessed ESWM on its core ability to predict missing elements of unseen transitions given a set of memories from an environment and a partially observable query. In both settings, we consider a closely related sequence-based model TEM-T as a baseline (Whittington et al., 2022). Instead of a set of disjoint memories, TEM-T receives a spanning trajectory as the memory bank and is queried to predict an unseen transition at the end of the trajectory. To enable fair comparison, the spanning trajectory is generated to be minimal (Extended Methods C.5).
We found a stark contrast between the model types in Open Arena setting (Fig. 2a). The ESWM-T model excelled at learning this task while ESWM-LSTM and ESWM-MAMBA models overfitted (see Fig. 8) and struggled to reach above chance on the validation set. This suggests that the attention mechanism in Transformer architectures—reminiscent of classic content-addressable memory models (Kanerva, 1988)—plays a key role in learning generalizable world models from episodic memories. However, simpler, hand-engineered memory-lookup mechanisms like those in Coda-Forno et al. (2022a) fail on this inference task (Fig. 17), highlighting the need for learnable attention-based mechanisms. Moreover, we also show that ESWM-T can also model spatial environments when trained on inputs distributions that lack the Minimality constraint (Fig. 9; see section D.2 for comparing ESWM-T trained with/without the Minimality constraint). In addition, the sequence-based transformer model TEM-T, while outperforming non-transformer ESWM models, still significantly underperforms relative to the transformer ESWM-T, highlighting the advantage of directly inferring from episodic memories, rather than segregated encoding of structure and memories. Furthermore, while ESWM-T clearly improves with larger model size, the performance in larger TEM-T models significantly degrades (Fig. 2a-b). For further discussions in this setting, see section D.1.
@@ 4.1 Modeling Spatial Environments from Episodic Memories
In the second scenario, the Random Wall, we found that all ESWM models are capable of learning the task (Fig. 2b). Although, there was a substantial gap between transformer-based models and other models, even with equal number of parameters (LSTM-4L, T-2L, and MAMBA-5L). This showed that ESWM learned a general strategy for modeling an environment from sparse and disjoint episodic memories, enabling it to generalize across environments with different structures. Consistent with our prediction, TEM-T fails at modeling the set of environments with varying structures. This underscores the importance of direct structural inference from memories rather than encoding it in the model’s weights.
Finally, we also considered environments with duplicate (non-unique) states. In the absence of historical context, one-step transition prediction prevents the model from disambiguating between duplicate states. However, this limitation can easily be overcome by prepending a trajectory, serving as context, to the query. To show this, we train ESWM-T in environments with duplicated states and find that it is capable of solving the task in both Open Arena and Random Wall (see Fig. 7).
@@
Figure 3: Spatial map emerges in ESWM’s latent space. a), b) ISOMAP projections of ESWM-T-2L’s activations for action prediction. Different columns are the same projection from different viewing angles. a) Spatial map in the absence of obstacles and b) in the presence of obstacles (a straight wall). c), d) ISOMAP projections of ESWM-T-14L’s activations. c) From left to right: A sample room with two disconnected regions whose shapes match boundaries’ shape; ESWM’s latent space when a memory bank observing either the top or bottom region is given as input; ESWM’s latent space when a memory bank observing both regions is given as input. d) From left to right: A sample room containing two disconnected regions whose shapes give no cues about their relative position; ESWM’s latent space when a memory bank of the room is given as input; Updated room with a new wall, the regions remain disconnected; ESWM’s latent space when a memory bank of the updated room is given as input. See Fig. 1014 for more examples.
4.2 Spatial Maps Emerge Robustly Within ESWM’s Latent Space
To simulate an unseen transition, ESWM must integrate disjoint memory fragments into a coherent internal map that reflects both observed states and obstacles. Inspired by low-dimensional, environment-like neural manifolds in rodents’ MTL Nakai et al. (2024), we applied ISOMAP Tenenbaum et al. (2000) (Extended Methods C.6) to ESWM’s latent space. The resulting manifolds revealed that ESWM consistently organizes its internal representations to mirror the spatial layout of each specific environment.
In obstacle-free environments, the manifold adopts a smooth, saddle-like shape (Fig. 3a). When obstacles are introduced, they create localized discontinuities that correspond precisely to blocked regions in physical space (Fig. 3b). This structure is remarkably robust and persists under out-of-distribution obstacle shapes (Fig. 10, 11), partial observation of environments (Fig. 12), across architectures (LSTM, Transformer, Mamba; Figs.1011), and throughout all transformer layers, which become progressively more continuous with depth (Fig. 14). All prediction tasks ($s_{s}$, $a$, $s_{e}$) yield similar maps (Fig. 12a). In the presence of disconnected memory clusters with unclear spatial relations, ESWM pieces its fragmented internal representation into a coherent map by aligning them along the inferred wall shapes or boundaries’ shapes (Fig. 3c,d; Additional Analysis D.3). Quantitatively, path lengths in latent and physical spaces correlate strongly ($R^{2}=0.89$, $N_{\text{path}}=1500$, $N_{\text{env}}=75$; Fig. 16b), and a linear classifier on ESWM activations can accurately discriminate which of two states is farther from a reference state with $93.42\pm 0.006\%$ accuracy (chance = 50%, $N_{\text{seed}}=10$; Extended Methods C.7). Furthermore, the sRSA (Levenstein et al., 2024) of a trained ESWM is significantly higher than that of an untrained one ($\text{trained}=0.78\pm 0.02$, $\text{untrained}=0.5\pm 0.04$, $n_{\text{env}}=1000$; paired $t$-test, tt=172.17, $p<10^{-3}$; Extended Methods C.8).
Notably, such structured spatial representations do not emerge in episodic agents trained for navigation, such as Episodic Planning Network (EPN; Ritter et al. (2020)), despite being trained with meta-RL on the same environments and data as ESWM (Fig. 10, 11; Appendix H).
4.3 ESWM Integrates Overlapping Memories
We ran several experiments to test whether ESWM truly retrieves the information scattered across its memory bank. First, we find that ESWM’s prediction uncertainty—measured by entropy of its output distribution probability—increases with the length of the integration path needed to solve a query (Fig. 4a). Second, we found that introducing shortcut memories that reduce this path length, significantly alters the prediction distribution (independent two-sample t-test; $n=5000$, $p<0.001$; Fig. 4b). Third, we find that ESWM trained in size-19 environments can generalize to substantially larger environments of size 37 in Open Arena, achieving lower, but still well-above-chance prediction accuracy ($s_{s}$: 59%, $a$: 90%, $s_{e}$: 55%)– unlikely if the model had simply memorized patterns. Fourth, ESWM’s performance improves with denser, non-minimal memory banks, despite being out-of-distribution in Random Wall (Fig. 4c). Lastly, despite being trained only on clean data, ESWM develops some robustness to noisy memory banks. Its prediction accuracy only dropped by 1% and 10%, respectively, when tested on memory banks containing contradictory state observations and transition outcomes (nn=5000; Extended Methods C.9). Altogether, these results suggest that ESWM indeed builds coherent spatial representations by integrating overlapping episodes.
(a)(b)(c)
@@ 4.3 ESWM Integrates Overlapping Memories
Figure 4: ESWM integrates memories. a) X-axis is the shortest path length between the source and end states in the query, with edges corresponding to episodic memories in the memory bank, and the Y-axis is the entropy of ESWM’s prediction probabilities (n=5000).The shaded region is 95% confidence interval computed over all predictions for each path length. b) KL divergence between ESWM prediction distributions before and after adding an episodic memory to the memory bank. An informative episodic memory shortens the integration path required for the model to solve the prediction task, while a non-informative episodic memory does not (n=5000). c) Prediction accuracy across memory bank sizes. In the gray region, memory banks minimally observe the environment; larger banks correspond to larger observed areas. In the blue region, the observed area is fixed (full environment), so larger memory banks correspond to a more densely observed area. ID and OOD denote in-distribution and out-of-distribution settings. (n=2000). ESWM-T-14L is used for a) and b) while ESWM-T-12L is used for c). All experiments are in Random Wall.
@@ 4.4 Exploration in Novel Environments
We designed a simple exploration algorithm based on ESWM’s “I don’t know” predictions (Alg.2). Starting with an empty memory bank, the agent evaluates all possible actions at each step by querying ESWM. It selects the action with the highest predicted uncertainty (i.e., highest “I don’t know” probability) and adds the resulting transition to memory. When ESWM is confident about all immediate actions, the agent performs multi-step lookahead by unrolling the model to identify frontier states with low prediction confidence (Yamauchi, 1997). A prediction is deemed unconfident if ESWM either outputs “I don’t know” or assigns $\leq 80\%$ probability to the predicted outcome.
Interestingly, the agent adopts an efficient zig-zag exploration strategy, only revisiting states when separated by long paths. In the Random Wall environment, ESWM explores 16.8% more unique states than EPN (one-way ANOVA; $n=1000$, $F=2803.63$, $p<0.05$; Tukey’s HSD post hoc test: ESWM >> EPN, $p<10^{-3}$), and achieves 96.48% of the performance of an oracle agent with full obstacle knowledge (Fig. 5a, Fig. 15). When exploration continues until memory reaches the max length seen during training, ESWM visits on average $91\pm 2$% of all unique states ($n_{trials}=1000$). Moreover, the collected memory banks allow ESWM to predict $s_{s}$, $a$, $s_{e}$ with 97%, 93%, 98% of the accuracy of those achieved using procedurally generated memory banks ($n=1000$). Together, these results highlight ESWM’s ability to support data-efficient, near-oracle exploration and mapping in novel environments.
@@
Figure 5: Exploration, navigation, and adaptability. a) Comparison of exploration strategies based on the number of unique states visited over 15 time steps in Random Wall. The optimal agent explores along the path found by the Traveling Salesman Algorithm over known free space. N=1000. b) Comparison of navigation success rate and path optimality over path lengths between EPN and ESWM, n=2400. c) Comparison of navigation success rates with increasing number of unexpected obstacles. The baseline agent is trained on the original environment and tested after structural changes, while ESWM navigates with memory banks from the original environment and needs to autonomously adapt to changes (n=100). The best seed for ESWM and EPN is used. We compare their performance across seeds in
Fig. 15. Random Wall Experiments include 19 locations.
@@ 4.5 Navigating with Minimal Spanning Episodic Memories
A key benefit of a world model is the ability to plan by simulating trajectories. We test this capability by using the memory-bank-equipped ESWM for planning in the Random Wall environment. The task is to find a path from a source state $s_{\text{source}}$ to a goal state $s_{\text{goal}}$ using only the current observation and the episodic memories in the memory bank, without access to global information like coordinates or a complete map of obstacles.
@@
Our planning strategy relies on ESWM’s ability to infer the end state of any given transition. Initialized at a state, an agent can imagine multiple steps into the future by querying ESWM on the consequences of all applicable actions, bootstrapping on ESWM’s own predicted end states, and tracking visited states until the goal state is reached (Alg.1). Then the agent can act out the imagined sequence of actions in reality.
@@
Using this strategy, an ESWM-based navigational agent can navigate between arbitrary states in Random Wall with 96.8% success rate and 99.2% path optimality, outperforming EPN Ritter et al. (2020) by 18% and 21% respectively (Fig. 5b) (one-sided independent t-test; n=2400: ESWM>>EPN on optimality; $p<10^{-3}$). Navigation is successful if the agent reaches the target within 15 time steps, and path optimality is defined as the ratio between the shortest and actual path length. We also include a variant of such an algorithm in the appendix (Additional Analysis D.4) that biasedly considers the action that shortens the geodesic distance to the target on ESWM’s latent spatial map, demonstrating how structured latent spaces in world models can serve as efficient heuristics for planning.
It is important to emphasize that the navigation success primarily stems from the high-fidelity world models constructed by ESWM using minimal memory (requiring $4\times$ fewer memories than EPN), rather than from the planning algorithm itself. Thus, the key finding is not that a classic search algorithm can outperform an RL agent, but that a world model learned from such sparse, disjoint data is accurate enough to enable it to do so.
Figure 6: ESWM in scaled-up environments a) (Top-Left) Examples of procedurally-generated MiniGrid environments. (Top-Right) Accuracies for $s_{s}$, $a$, and $s_{e}$ predictions for ESWM-T-12L. (Bottom) ISOMAP analysis on ESWM’s latent space shows a map-like representation. Inset shows the color scheme across spatial locations. b) (Top-Left) Examples of procedurally-generated ProcThor environments. (Top-Right) A sample transition. States are high-dimensional images, and actions are continuous 3-dimensional vectors. (Bottom) ESWM-T-12L’s performance in unseen test environments for action (measured by accuracy on sub-components $\Delta xy$ and $\Delta\theta$) and end-state prediction (measured by cosine similarity to the frozen vision model’s representation).